OpenStudy (anonymous):

Neil skated 7/8 mile. Zavier skated 3/4 mile. How much farther did Neil skate than Zavier?

Mehek (mehek14):

\( \dfrac{7}{8} = 0.875 \\ \dfrac{3}{4} = 0.75\\ 0.875 - 0.75 = 0.125\\ 0.125 = \dfrac{1}{8}\)

Mehek (mehek14):

or if you want to subtract by just fractions \(\large \dfrac{7}{8} - \dfrac{6}{8} = \dfrac{1}{8}\)

OpenStudy (anonymous):

ok